# Asymptotically Unbiased Synthetic Control Methods by Density Matching

### v4

@@ Abstract
Synthetic Control Methods (SCMs) have become a fundamental tool for comparative case studies. The core idea behind SCMs is to estimate treatment effects by predicting counterfactual outcomes for a treated unit using a weighted combination of observed outcomes from untreated units. The accuracy of these predictions is crucial for evaluating the treatment effect of a policy intervention. Subsequent research has therefore focused on estimating SC weights. In this study, we highlight a key endogeneity issue in existing SCMs—namely, the correlation between the outcomes of untreated units and the error term of the synthetic control, which leads to bias in both counterfactual outcome prediction and treatment effect estimation. To address this issue, we propose a novel SCM based on density matching, assuming that the outcome density of the treated unit can be approximated by a weighted mixture of the joint density of untreated units. Under this assumption, we estimate SC weights by matching the moments of the treated outcomes with the weighted sum of the moments of the untreated outcomes. Our method offers three advantages: first, under the mixture model assumption, our estimator is asymptotically unbiased; second, this asymptotic unbiasedness reduces the mean squared error in counterfactual predictions; and third, our method provides full densities of the treatment effect rather than just expected values, thereby broadening the applicability of SCMs. Finally, we present experimental results that demonstrate the effectiveness of our approach. [^1] This study was presented at the Workshop on Counterfactuals in Minds and Machines at the International Conference on Machine Learning in July 2023 and at the International Conference on Econometrics and Statistics in August 2023.
@@ 1 Introduction
Synthetic control methods (SCMs, Abadie and Gardeazabal, 2003; Abadie et al., 2010) have emerged as a crucial tool for program evaluation in comparative case studies across disciplines such as economics, statistics, and machine learning. As noted by Athey and Imbens (2017), SCMs represent “arguably the most important innovation in the policy evaluation literature in the last 15 years” due to their “simplicity of the idea, and the obvious improvement over standard methods.”
@@
In SCMs, there are multiple units, where one unit receives a policy intervention (treated unit) at a certain time while the remaining units remain untreated. The objective is to estimate the treatment effect for the treated unit, defined as the difference between its observed factual outcome and its unobserved counterfactual outcome. Untreated units help account for unobserved trends in the outcome over time that are unrelated to the policy intervention. The key idea is that an optimally weighted combination of the untreated units, referred to as the synthetic control (SC) unit, serves as an appropriate estimate of the counterfactual outcome. See Figure 1 for an illustration. d SCMs have been widely applied in various empirical studies, such as evaluating the impact of terrorism on GDP (Abadie and Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham and Shah, 2017). For instance, Abadie et al. (2010) applies SCM to study the treatment effect of a comprehensive tobacco control program implemented in California in 1988. While smoking rates in California declined following the program’s implementation, it remained unclear whether this decrease was attributable to the policy. For estimating the treatment effect, annual per-capita cigarette sales data from states that did not implement a similar tobacco control program were used as untreated units. SCM was then applied to construct California’s counterfactual outcome, representing the smoking rate in the absence of the tobacco control program.
@@
A key challenge in SCMs is the assumption, made in most existing studies (Abadie, 2002), that the expected outcome of the treated unit can be represented as a weighted sum of the outcomes of untreated units. Ferman and Pinto (2021) highlights that this assumption leads to a non-negligible bias in the estimator. According to Ferman and Pinto (2021), this bias arises from endogeneity in the linear regression model implied by the SCM assumption due to the correlation between the outcomes of untreated units and the error term. This issue is analogous to the endogeneity caused by measurement error (Greene, 2003). We refer to this problem as implicit endogeneity.
@@
To address this issue, we propose an alternative approach based on a mixture model assumption rather than a linear relationship among expected outcomes. Specifically, we assume that the density of the treated unit’s outcome can be approximated by a weighted combination of the densities of the untreated units. This assumption further implies that each moment of the treated unit’s outcome can be represented as a weighted sum of the moments of the untreated units. Leveraging this insight, we estimate SC weights by matching the moments of the treated unit’s outcome to those of the untreated units. We refer to our method as the Density Matching SCM (DMSCM). Unlike existing methods, our proposed estimator is asymptotically unbiased if the mixture model assumption is correct. We validate the effectiveness of our approach through both simulation studies and empirical analysis.
@@
The mixture model assumption has appeared in previous and concurrent studies, including Wan et al. (2018), Gunsilius (2023), and Shi et al. (2022). Our proposed method can be viewed as a simplification of the distributional SCMs introduced by Gunsilius (2023). While Gunsilius (2023) does not explicitly address the implicit endogeneity problem, they note that their results diverge from those obtained under standard SCMs. This discrepancy arises because, under the mixture model assumption, Gunsilius (2023) provides a consistent estimator for the treatment effect, whereas existing SCM methods do not. We establish a connection between distributional SCMs and the asymptotic bias problem and propose a simplified estimation method based on density matching.
@@
  • Proposing a density-matching approach for asymptotically unbiased SCMs.
  • Simplifying the implementation of density matching in SCMs.
We formulate the problem in Section 2. Next, in Section 3, we review the basic idea of SCMs. In Section 4, we highlight the problem of implicit endogeneity. In Section 5, we then propose our method under the mixture model assumption. We investigate the asymptotic properties of this method in Section 6 and present a statistical inference method in Section 7. In Section 8, we discuss related topics and existing studies. Finally, in Section 9, we examine the performance of our estimators through extensive simulations.
@@
Related work. Following the pioneering work by Abadie and Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015). These SCMs, including their variants such as those by Doudchenko and Imbens (2016) and Ben-Michael et al. (2019), are based on minimizing the pre-treatment sum of squared errors between the outcome of the treated unit and the weighted sum of those of the untreated units, with some regularization and constraints regarding the weights. We refer to such SCMs as Least-Squares SCMs (LS-SCMs).
@@
Abadie et al. (2010) demonstrates that under a pre-treatment fit, the bias of the LS-SCM estimator converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases. However, Ferman and Pinto (2021) finds that if the pre-treatment fit does not hold, then the LS-SCM estimator has a bias and is not consistent. The cause of this bias is the correlation between the outcomes of the untreated units and the error term in the posited linear model, which we refer to as implicit endogeneity.
Several methods have been proposed to address this bias issue. Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where both the number of untreated units ($J\to\infty$) and the length of the pre-intervention periods ($T_{0}$) grow large. Concurrently with us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias reported by Ferman and Pinto (2021). [^2] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and he and we conducted our studies completely independently. While our formulation and his regarding endogeneity are similar, our approach differs significantly. We address the problem through density matching, similar to the strategy in Gunsilius (2023), whereas Fry (2024) employs an instrumental variable approach using outcomes of untreated units not used in SCMs as instruments.
@@
This study also addresses a question raised by Gunsilius (2023). While Gunsilius (2023) suggests that standard least squares SCMs try to match the first moment of the outcome distribution (and therefore differ from distributional approaches), we stress that least squares approaches are additionally vulnerable to implicit endogeneity, which can generate bias. Our density-matching approach avoids this problem by matching a broader set of distributional features.
@@
Ours DMSCM $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\ \ \ \forall y\in\mathbb{R}$ Unbiased
Abadie and Gardeazabal (2003) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Biased
@@ 2 Problem Setting
Suppose there are $J+1$ units, indexed by $j\in\mathcal{J}\coloneqq\{0,1,2,\dots,J\}$, and a time series indexed by $t\in\mathcal{T}=\{1,2,\dots,T\}$, where $3\leq T\in\mathbb{N}$. Without loss of generality, assume that the first unit ($j=0$) is the treated unit, i.e., the unit affected by the policy intervention of interest. The set of potential control units is $\mathcal{J}^{U}\coloneqq\mathcal{J}\backslash\{0\}=\{1,2,\dots,J\}$, consisting of units not affected by the intervention. We also assume that our data span $T$ periods and that the intervention occurs at $t=T_{0}$, with $2\leq T_{0}<T$. Observations in period $t\in\mathcal{T}_{0}\coloneqq\{1,2,\dots,T_{0}\}$ are from before the intervention, and observations in period $t\in\mathcal{T}_{1}\coloneqq\mathcal{T}\backslash\mathcal{T}_{0}$ are from after the intervention. Let $T_{1}\coloneqq|\mathcal{T}_{1}|=T-T_{0}$.
@@ 2.3 Treatment Effect
Distributional Treatment Effects. In program evaluation, we are often interested not only in the scalar treatment effect but also in the distributions of outcomes (Maier, 2011) or their functionals, such as quantile treatment effects (QTEs) and expected social welfare (Abadie, 2002). By applying our method, we can also estimate the distribution of the counterfactual outcome, $p^{N}_{0,t}(y)$, for $t\in\mathcal{T}_{1}$. The treatment effect is then captured by the difference between $p^{I}_{0,t}(y)$ (the distribution of the actual outcome) and $p^{N}_{0,t}(y)$, or by functionals of these distributions, such as QTEs and expected social welfare. Depending on the application, we may also adjust the target density, for example, by averaging across all post-intervention periods: $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$. We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus and Oprescu, 2022; Chikahara et al., 2022). In the context of SCMs, Gunsilius (2023) and Chen (2020) have proposed methods for estimating such distributional treatment effects.
@@ 3 Recap of SCMs
This section reviews the original SCM proposed by Abadie and Gardeazabal (2003). To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ using a weighted sum of $\{Y^{N}_{j,t}\}_{j=1}^{J}$ for $t\in\mathcal{T}_{1}$. Specifically, SCMs predict $Y^{N}_{0,t}$ as
@@
Several methods exist for estimating the SC weights. In the SCM proposed by Abadie and Gardeazabal (2003), these weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. Define
@@
Then, Abadie and Gardeazabal (2003) estimate the SC weights as
@@
The literature on inference for SCMs has grown significantly, extending the original proposals by Abadie and Gardeazabal (2003) and Abadie et al. (2010). Notable contributions include works by Li (2020), Shaikh and Toulis (2021), Cattaneo et al. (2021), and Chernozhukov et al. (2019). Doudchenko and Imbens (2016) proposes estimating the SC weight as
@@ 4.1 Linear Models in Expected Outcomes
This assumption, often called the perfect pre-treatment fit (Ferman and Pinto, 2021), implies that the stochastic process $\left(Y^{N}_{0,t}\right)_{t\in\mathcal{T}_{0}}$ can be fully replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$ for $t\in\mathcal{T}_{0}$. In linear factor models where $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is often considered unrealistic. Note that this is a specific case and not necessary for $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$.
@@ 4.2 Asymptotic Bias in Linear Factor Models
Under these assumptions, Ferman and Pinto (2021) shows that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{0,t}$ suffers from bias and is not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero. We also adopt the following assumption regarding the linear factor model.
@@
Under Assumptions 4.2 and 4.3, Ferman and Pinto (2021) concludes that the LS-SCM proposed by Abadie and Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$, as the following proposition.
Proposition 4.4 (Proposition 1 in Ferman and Pinto (2021)).
@@
To alleviate this bias problem, Ferman and Pinto (2021) proposes the demeaned SC estimator:
@@ 4.3 Implicit Endogeneity
We now interpret the asymptotic bias found by Ferman and Pinto (2021) as an endogeneity problem. We consider (2), that is, $\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}]$, which is more general than the linear factor model.
@@
Thus, the identification of treatment effects faces a challenge due to implicit endogeneity. While the estimator $\bm{w}^{\mathrm{Abadie}}$ remains standard in many applications, it inherits this endogeneity issue. Such concerns have been noted in prior works (Chernozhukov et al., 2019; Ferman and Pinto, 2021; Ferman, 2021), and Fry (2024) independently frames the bias as an endogeneity problem, proposing an instrumental variable method to address it.
5 SCM by Density Matching
As discussed in Section 4, Assumption 4.1 can be weak or incompatible for identifying $\bm{w}^{*}$ when using least squares. We, therefore, adopt a different assumption for identification. Specifically, we propose that the density of $Y^{N}_{0,t}$ can be replicated by a weighted combination of the densities of $Y^{N}_{j,t}$, and we introduce an SCM based on density matching.

### v5

# Asymptotically Unbiased Synthetic Control Methods by Moment Matching

@@ Abstract
Synthetic Control Methods (SCMs) have become a fundamental tool for comparative case studies. The core idea behind SCMs is to estimate treatment effects by predicting counterfactual outcomes for a treated unit using a weighted combination of observed outcomes from untreated units. The accuracy of these predictions is crucial for evaluating the treatment effect of a policy intervention. Subsequent research has therefore focused on estimating SC weights. In this study, we highlight a key endogeneity issue in existing SCMs—namely, the correlation between the outcomes of untreated units and the error term of the synthetic control, which leads to bias in both counterfactual outcome prediction and treatment effect estimation. To address this issue, we propose a novel SCM based on moment matching, assuming that the outcome distribution of the treated unit can be approximated by a weighted mixture of the distributions of untreated units. Under this assumption, we estimate SC weights by matching the moments of the treated outcomes with the weighted sum of the moments of the untreated outcomes. Our method offers three advantages: first, under the mixture model assumption, our estimator is asymptotically unbiased; second, this asymptotic unbiasedness reduces the mean squared error in counterfactual predictions; and third, our method provides full distributions of the treatment effect rather than just expected values, thereby broadening the applicability of SCMs. Finally, we present experimental results that demonstrate the effectiveness of our approach.
@@ 1 Introduction
Synthetic control methods (SCMs, Abadie & Gardeazabal, 2003; Abadie et al., 2010) have emerged as a crucial tool for program evaluation in comparative case studies across disciplines such as economics, statistics, and machine learning. In SCMs, there are multiple units, where one unit receives a policy intervention (treated unit) at a certain time while the remaining unitts remain untreated. The objective is to estimate the treatment effect for the treated unit, defined as the difference between its observed factual outcome and its unobserved counterfactual outcome. Untreated units help account for unobserved trends in the outcome over time that are unrelated to the policy intervention. The key idea is that an optimally weighted combination of the untreated units, referred to as the synthetic control (SC) unit, serves as an appropriate estimate of the counterfactual outcome. See Figure 1 for an illustration.
SCMs have been widely applied in various empirical studies, such as evaluating the impact of terrorism on GDP (Abadie & Gardeazabal, 2003) and the decriminalization of indoor prostitution (Cunningham & Shah, 2017). For instance, Abadie et al. (2010) applies SCM to study the treatment effect of a comprehensive tobacco control program implemented in California in 1988. While smoking rates in California declined following the program’s implementation, it remained unclear whether this decrease was attributable to the policy. For estimating the treatment effect, annual per-capita cigarette sales data from states that did not implement a similar tobacco control program were used as untreated units. SCM was then applied to construct California’s counterfactual outcome, representing the smoking rate in the absence of the tobacco control program.
@@
A key challenge in SCMs is the assumption, made in most existing studies (Abadie, 2002), that the expected outcome of the treated unit can be represented as a weighted sum of the outcomes of untreated units. Ferman & Pinto (2021) highlights that this assumption leads to a non-negligible bias in the estimator. According to Ferman & Pinto (2021), this bias arises from endogeneity in the linear regression model implied by the SCM assumption due to the correlation between the outcomes of untreated units and the error term. This issue is analogous to the endogeneity caused by measurement error (Greene, 2003). We refer to this problem as implicit endogeneity.
@@
To address this issue, we propose an alternative approach based on a mixture model assumption rather than a linear relationship among expected outcomes. Specifically, we assume that the distribution of the treated unit’s outcome can be approximated by a weighted combination of the distributions of the untreated units. This assumption further implies that each moment of the treated unit’s outcome can be represented as a weighted sum of the moments of the untreated units. Leveraging this insight, we estimate SC weights by matching the moments of the treated unit’s outcome to those of the untreated units. We refer to our method as the Moment Matching SCM (MMSCM). Unlike existing methods, our proposed estimator is asymptotically unbiased if the mixture model assumption is correct. We validate the effectiveness of our approach through both simulation studies and empirical analysis.
@@
The mixture model assumption has appeared in previous and concurrent studies, including Wan et al. (2018), Gunsilius (2023), and Shi et al. (2022). Our proposed method can be viewed as a simplification of the distributional SCMs introduced by Gunsilius (2023). While Gunsilius (2023) does not explicitly address the implicit endogeneity problem, they note that their results diverge from those obtained under standard SCMs. This discrepancy arises because, under the mixture model assumption, Gunsilius (2023) provides a consistent estimator for the treatment effect, whereas existing SCM methods do not. We establish a connection between distributional SCMs and the asymptotic bias problem and propose a simplified estimation method based on moment matching.
@@
  • Proposing a moment-matching approach for asymptotically unbiased SCMs.
  • Simplifying the implementation of moment matching in SCMs.
We formulate the problem in Section 2. Next, in Section 3, we review the basic idea of SCMs. In Section 4, we highlight the problem of implicit endogeneity. In Section 5, we then propose our method under the mixture model assumption. We investigate the asymptotic properties of this method in Section 5.5 and present a statistical inference method in Section 5.6. In Section 6, we discuss related topics and existing studies. Finally, in Section 7, we examine the performance of our estimators through extensive simulations.
@@
Related work. Following the pioneering work by Abadie & Gardeazabal (2003), numerous studies have advanced the use of SCMs in comparative studies (Abadie et al., 2010, 2015). These SCMs, including their variants such as those by Doudchenko & Imbens (2016) and Ben-Michael et al. (2019), are based on minimizing the pre-treatment sum of squared errors between the outcome of the treated unit and the weighted sum of those of the untreated units, with some regularization and constraints regarding the weights. We refer to such SCMs as Least-Squares SCMs (LS-SCMs).
@@
Abadie et al. (2010) demonstrates that under a pre-treatment fit, the bias of the LS-SCM estimator converges to zero asymptotically as the length of the pre-treatment periods $T_{0}$ increases. However, Ferman & Pinto (2021) finds that if the pre-treatment fit does not hold, then the LS-SCM estimator has a bias and is not consistent. The cause of this bias is the correlation between the outcomes of the untreated units and the error term in the posited linear model, which we refer to as implicit endogeneity.
Several methods have been proposed to address this bias issue. Ferman (2021) demonstrates that we can obtain an asymptotically unbiased estimator under the regime where both the number of untreated units ($J\to\infty$) and the length of the pre-intervention periods ($T_{0}$) grow large. Concurrently with us, Fry (2024) also discusses the endogeneity problem in SCMs as a cause of the asymptotic bias reported by Ferman & Pinto (2021). [^1] Note that his study became public on arXiv after our study became public on arXiv and at the ICML workshop, and he and we conducted our studies completely independently. While our formulation and his regarding endogeneity are similar, our approach differs significantly. We address the problem through moment matching, similar to the strategy in Gunsilius (2023), whereas Fry (2024) employs an instrumental variable approach using outcomes of untreated units not used in SCMs as instruments.
@@
This study also addresses a question raised by Gunsilius (2023). While Gunsilius (2023) suggests that standard least squares SCMs try to match the first moment of the outcome distribution (and therefore differ from distributional approaches), we stress that least squares approaches are additionally vulnerable to implicit endogeneity, which can generate bias. Our moment-matching approach avoids this problem by matching a broader set of distributional features.
@@
Ours MMSCM $F_{Y^{N}_{0,t}}(y)=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}F_{Y^{N}_{j,t}}(y)\ \ \ \forall y\in\mathbb{R}$ Unbiased
Abadie & Gardeazabal (2003) Constrained least squares $\mathbb{E}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}[Y^{N}_{j,t}]$ Biased
@@ 2 Problem Setting
[…]
@@ 2.3 Treatment Effect
Distributional Treatment Effects. In program evaluation, we are often interested not only in the scalar treatment effect but also in the distributions of outcomes (Maier, 2011) or their functionals, such as quantile treatment effects (QTEs) and expected social welfare (Abadie, 2002). By applying our method, we can also estimate the distribution of the counterfactual outcome, $p^{N}_{0,t}(y)$, for $t\in\mathcal{T}_{1}$. The treatment effect is then captured by the difference between $p^{I}_{0,t}(y)$ (the distribution of the actual outcome) and $p^{N}_{0,t}(y)$, or by functionals of these distributions, such as QTEs and expected social welfare. Depending on the application, we may also adjust the target density, for example, by averaging across all post-intervention periods: $\overline{p}^{N}_{0}(y)=\frac{1}{T_{1}}\sum_{t\in\mathcal{T}_{1}}p^{N}_{0,t}(y)$. We refer to the treatment effects related to these distributions as distributional treatment effects (Park et al., 2021; Kallus & Oprescu, 2022; Chikahara et al., 2022). In the context of SCMs, Gunsilius (2023) and Chen (2020) have proposed methods for estimating such distributional treatment effects.
@@ 3 Recap of SCMs
This section reviews the original SCM proposed by Abadie & Gardeazabal (2003). To estimate the treatment effect $\tau_{0,t}$, we predict the counterfactual outcome $Y^{N}_{0,t}$ using a weighted sum of $\{Y^{N}_{j,t}\}_{j=1}^{J}$ for $t\in\mathcal{T}_{1}$. Specifically, SCMs predict $Y^{N}_{0,t}$ as
@@
Several methods exist for estimating the SC weights. In the SCM proposed by Abadie & Gardeazabal (2003), these weights are determined by minimizing the squared error between $Y^{N}_{0,t}$ and the weighted sum of $Y^{N}_{1,t},\dots,Y^{N}_{J,t}$. Define
@@
Then, Abadie & Gardeazabal (2003) estimate the SC weights as
@@
The literature on inference for SCMs has grown significantly, extending the original proposals by Abadie & Gardeazabal (2003) and Abadie et al. (2010). Notable contributions include works by Li (2020), Shaikh & Toulis (2021), Cattaneo et al. (2021), and Chernozhukov et al. (2021). Doudchenko & Imbens (2016) proposes estimating the SC weight as
@@ 4.1 Linear Models in Expected Outcomes
This assumption, often called the perfect pre-treatment fit (Ferman & Pinto, 2021), implies that the stochastic process $\left(Y^{N}_{0,t}\right)_{t\in\mathcal{T}_{0}}$ can be fully replicated by $\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$ for $t\in\mathcal{T}_{0}$. In linear factor models where $Y^{N}_{j,t}=c_{j}+\delta_{t}+\lambda_{t}\mu_{j}+\epsilon_{j,t}$, this assumption might imply $\epsilon_{0,t}=\sum_{j\in[J]}w^{*}_{j}\epsilon_{j,t}$, which is often considered unrealistic. Note that this is a specific case and not necessary for $Y^{N}_{0,t}=\sum_{j\in[J]}w^{*}_{j}Y^{N}_{j,t}$.
@@ 4.2 Asymptotic Bias in Linear Factor Models
Under these assumptions, Ferman & Pinto (2021) shows that the SC estimator $\widehat{\tau}^{\mathrm{Abadie}}_{0,t}$ suffers from bias and is not consistent for $\tau_{0,t}$ when $c_{j}$ is non-zero. We also adopt the following assumption regarding the linear factor model.
@@
Under Assumptions 4.2 and 4.3, Ferman & Pinto (2021) concludes that the LS-SCM proposed by Abadie & Gardeazabal (2003) does not yield a consistent estimator for $\tau_{0,t}$, as the following proposition.
Proposition 4.4 (Proposition 1 in Ferman & Pinto (2021)).
@@
To alleviate this bias problem, Ferman & Pinto (2021) proposes the demeaned SC estimator:
@@ 4.3 Implicit Endogeneity
We now interpret the asymptotic bias found by Ferman & Pinto (2021) as an endogeneity problem. We consider (2), that is, $\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}]$, which is more general than the linear factor model.
@@
Thus, the identification of treatment effects faces a challenge due to implicit endogeneity. While the estimator $\widehat{\bm{w}}^{\mathrm{Abadie}}_{n}$ remains standard in many applications, it inherits this endogeneity issue. Such concerns have been noted in prior works (Chernozhukov et al., 2021; Ferman & Pinto, 2021; Ferman, 2021), and Fry (2024) independently frames the bias as an endogeneity problem, proposing an instrumental variable method to address it.
Bias is caused by a measurement error. As we showed above, the source of the implicit endogeneity is the measurement error. Since we assume $\mathbb{E}_{0,t}[Y^{N}_{0,t}]=\sum_{j\in\mathcal{J}^{U}}w^{*}_{j}\mathbb{E}_{j,t}[Y^{N}_{j,t}]$ as a regression model, the regressors are $(\mathbb{E}_{j,t}[Y^{N}_{j,t}])_{j\in\mathcal{J}^{U}}$. However, in practice, we can only use $(Y_{j,t})$ as the regressors. Because the regressors contain measurement errors $(\mathbb{E}_{j,t}[Y^{N}_{j,t}]-Y_{j,t})_{j\in\mathcal{J}^{U}}$, endogeneity arises, as is commonly discussed in elementary textbooks (Greene, 2003).
5 SCM by Moment Matching
As discussed in Section 4, Assumption 4.1 can be weak or incompatible for identifying $\bm{w}^{*}$ when using least squares. We, therefore, adopt a different assumption for identification. Specifically, we propose that the distribution of $Y^{N}_{0,t}$ can be replicated by a weighted combination of the distributions of $Y^{N}_{j,t}$, and we introduce an SCM based on moment matching.